# Reference Points, Risk-Taking Behavior, and Competitive Outcomes in Sequential Settings

### v3

@@ 8 Conclusion
Counterfactual simulations show that without competitive pressure, lifters tend to adopt more conservative attempts. However, the effects on performance are highly heterogeneous: while some lifters benefit from the absence of pressure, some lifters experience lower success rates and reduced expected outcomes. These results highlight that external competition plays a dual role—motivating some lifters to perform better while destabilizing others—underscoring the importance of individual sensitivity in shaping performance under pressure.
@@
Appendix A Online Appendix (Not for publication)
A.1 Comprehensive Attempt Stage Model
In this section, we present an attempt stage model that simultaneously considers the influence of both the lower and higher rivals on the lifter’s weight declaration. By accounting for both sources of competitive pressure simultaneously, this model provides a more complete representation of how lifters adjust their declared weights in response to the surrounding competition. However, the added complexity makes the model less intuitive to interpret, as the interactions between the two rivals’ effects become more intricate.
We maintain the notation used in Section 3, let $w_{L}$ denote the lower rival’s actual attempt weight which is observable to lifter $i$ during the attempt stage. Also, let $w_{H}$ denote the higher rival’s expected attempt weight under lifter $i$’s rational expectations. Although this is not observable to lifter $i$’s attempt stage, we treat this as highly predictable information to lifter $i$. Additionally, let $y_{L}$, $y_{H}$, and $y_{i}$ represent the best outcome weights from earlier attempts within the competition for the lower rival, the higher rival, and lifter $i$, respectively, which are known in lifter $i$’s current attempt stage. The (strict) inequality conditions $y_{L}<y_{i}<y_{H}$, determined by rank order, and $y_{L}<w_{L}$, $y_{i}<w_{i}$, and $y_{H}<w_{H}$, meaning that the outcome weight must not exceed the corresponding attempt weight, must hold. Although we can assume that $w_{L}<w_{H}$, meaning that the lower rival does not attempt to overtake two higher-ranked rivals as inferred from our data, must hold, we keep generality and allow $w_{L}>w_{H}$. These conditions determine the joint probability of the outcome tuple as shown in Table 13.
Table 13: Joint probability of outcome tuple
Pressure $i$ $L$ $H$ Outcome $i$’s rank Probability
$w_{i}<\min\{w_{L},w_{H}\}$ S S S $w_{i}<\min\{w_{L},w_{H}\}$- $q(w)q(w_{L})q(w_{H})$
$w_{i}<\min\{w_{L},w_{H}\}$ S S F $w_{i}<\min\{w_{L},y_{H}\}$- $q(w)q(w_{L})(1-q(w_{H}))P(w_{i}<y_{H})$
$w_{i}<\min\{w_{L},w_{H}\}$ S S F $y_{H}<w_{i}<w_{L}$ K $q(w)q(w_{L})(1-q(w_{H}))P(w_{i}>y_{H})$
$w_{i}<\min\{w_{L},w_{H}\}$ S F S $y_{L}(<y_{i})<w_{i}<w_{H}$ K $q(w)(1-q(w_{L}))q(w_{H})$
$w_{i}<\min\{w_{L},w_{H}\}$ S F F $y_{L}(<y_{i})<w_{i}<y_{H}$ K $q(w)(1-q(w_{L}))(1-q(w_{H}))P(w_{i}<y_{H})$
$w_{i}<\min\{w_{L},w_{H}\}$ S F F $y_{L}<y_{H}<w_{i}$ + $q(w)(1-q(w_{L}))(1-q(w_{H}))P(w_{i}>y_{H})$
$w_{i}<\min\{w_{L},w_{H}\}$ F S Any $y_{i}<\min\{w_{L},y_{H}\}$- $(1-q(w))q(w_{L})$
$w_{i}<\min\{w_{L},w_{H}\}$ F F Any $y_{L}<y_{i}<y_{H}$ K $(1-q(w))(1-q(w_{L}))$
$w_{L}<w_{i}<w_{H}$ S S S $w_{L}<w_{i}<w_{H}$ K $q(w)q(w_{L})q(w_{H})$
$w_{L}<w_{i}<w_{H}$ S S F $w_{L}<w_{i}<y_{H}(<w_{H})$ K $q(w)q(w_{L})(1-q(w_{H}))P(w_{i}<y_{H})$
$w_{L}<w_{i}<w_{H}$ S S F $\max\{w_{L},y_{H}\}<w_{i}$ + $q(w)q(w_{L})(1-q(w_{H}))P(w_{i}>y_{H})$
$w_{L}<w_{i}<w_{H}$ S F S $y_{L}(<w_{L})<w_{i}<w_{H}$ K $q(w)(1-q(w_{L}))q(w_{H})$
$w_{L}<w_{i}<w_{H}$ S F F $y_{L}<w_{i}<y_{H}(<w_{H})$ K $q(w)(1-q(w_{L}))(1-q(w_{H}))P(w_{i}<y_{H})$
$w_{L}<w_{i}<w_{H}$ S F F $y_{L}<y_{H}<w_{i}$ + $q(w)(1-q(w_{L}))(1-q(w_{H}))P(w_{i}>y_{H})$
$w_{L}<w_{i}<w_{H}$ F S S $y_{i}<w_{L}<w_{H}$- $(1-q(w))q(w_{L})q(w_{H})P(w_{L}>y_{i})$
$w_{L}<w_{i}<w_{H}$ F S S $w_{L}<y_{i}<w_{H}$ K $(1-q(w))q(w_{L})q(w_{H})P(w_{L}<y_{i})$
$w_{L}<w_{i}<w_{H}$ F S F $w_{L}<y_{i}<y_{H}$ K $(1-q(w))q(w_{L})(1-q(w_{H}))P(w_{L}<y_{i})$
$w_{L}<w_{i}<w_{H}$ F S F $y_{i}<\min\{w_{L},y_{H}\}$- $(1-q(w))q(w_{L})(1-q(w_{H}))P(w_{L}>y_{i})$
$w_{L}<w_{i}<w_{H}$ F F S $y_{L}<y_{i}<w_{H}$ K $(1-q(w))(1-q(w_{L}))q(w_{H})$
$w_{L}<w_{i}<w_{H}$ F F F $y_{L}<y_{i}<y_{H}$ K $(1-q(w))(1-q(w_{L}))(1-q(w_{H}))$
$\max\{w_{L},w_{H}\}<w_{i}$ S Any Any $\max\{w_{L},w_{H}\}<w_{i}$ + $q(w)$
$\max\{w_{L},w_{H}\}<w_{i}$ F S Any $w_{L}<y_{i}<\min\{y_{H},w_{H}\}$ K $(1-q(w))q(w_{L})P(w_{L}<y_{i})$
$\max\{w_{L},w_{H}\}<w_{i}$ F S Any $y_{i}<\min\{w_{L},y_{H}\}$- $(1-q(w))q(w_{L})P(w_{L}>y_{i})$
$\max\{w_{L},w_{H}\}<w_{i}$ F F Any $y_{L}<y_{i}<\min\{y_{H},w_{H}\}$ K $(1-q(w))(1-q(w_{L}))$
Notes: F = failed, S = success, K = keep. We ignore the equality case like $w_{L}=y_{i}$ in the probability because the body weight information works as a tie-breaker for the case.
Given the values of $w_{L},w_{H},y_{L},y_{H}$, and $y_{i}$ with the rank-dependent value function in Section 3, the expected evaluation of the lottery $(w_{i},q(w_{i}))$ for lifter $i$ with reference point $r_{i}$ is
$EU_{i}=v(x_{i}-r_{i})q(w_{i})$
=$=\left\{\begin{array}[]{ll}\Bigl{(}q(w_{i})q(w_{L})q(w_{H})+q(w_{i})q(w_{L})(1-q(w_{H}))P(w_{i}<y_{H})+(1-q(w_{i}))q(w_{L})\Bigr{)}[-\lambda]&\text{if }w_{i}<w_{L},w_{H}\\ +\Bigl{(}q(w_{i})(1-q(w_{L}))(1-q(w_{H}))P(w_{i}>y_{H})\Bigr{)}[1]\\ \Bigl{(}(1-q(w_{i}))q(w_{L})q(w_{H})P(w_{L}>y_{i})+(1-q(w_{i}))q(w_{L})(1-q(w_{H}))P(w_{L}>y_{i})\Bigr{)}[-\lambda]&\text{if }w_{L}<w_{i}<w_{H}\\ +\Bigl{(}q(w_{i})q(w_{L})(1-q(w_{H}))P(w_{i}>y_{H})+q(w_{i})(1-q(w_{L}))(1-q(w_{H}))P(w_{i}>y_{H})\Bigr{)}[1]\\ \Bigl{(}(1-q(w_{i}))q(w_{L})P(w_{L}>y_{i})\Bigr{)}[-\lambda]+q(w_{i})[1]&\text{if }w_{L},w_{H}<w_{i}\end{array}\right.$ (8)
=$=\left\{\begin{array}[]{ll}\Bigl{(}q(w_{i})q(w_{L})q(w_{H})+q(w_{i})q(w_{L})(1-q(w_{H}))P(w_{i}<y_{H})+(1-q(w_{i}))q(w_{L})\Bigr{)}[-\lambda]&\text{if }w_{i}<w_{L},w_{H}\\ +\Bigl{(}q(w_{i})(1-q(w_{L}))(1-q(w_{H}))P(w_{i}>y_{H})\Bigr{)}[1]\\ \Bigl{(}(1-q(w_{i}))q(w_{L})P(w_{L}>y_{i})\Bigr{)}[-\lambda]+\Bigl{(}q(w_{i})(1-q(w_{H}))P(w_{i}>y_{H})\Bigr{)}[1]&\text{if }w_{L}<w_{i}<w_{H}\\ \Bigl{(}(1-q(w_{i}))q(w_{L})P(w_{L}>y_{i})\Bigr{)}[-\lambda]+q(w_{i})[1]&\text{if }w_{L},w_{H}<w_{i}\end{array}\right.$ (13)
Although the representation of the expected evaluation of the lottery in Equation (13) may appear complex, in particular, when $w_{L}<w_{i}<w_{H}$, it reveals that there are discontinuities at $w_{L}$ and $w_{H}$. Additionally, it is crucial to note that $P(w_{i}>y_{H})$ (that is, chance of the higher rival’s being turned around by the lifter) and $P(w_{L}>y_{i})$ (that is, chance of the lower rival’s turning around the lifter) are key factors that capture the pressure exerted by lower and higher rivals, respectively.
Even without further specification, Equation (13) offers theoretical predictions regarding the optimal choice of attempt weight $w_{i}$ for each domain of $w_{i}$. First, if lifter $i$ aims to choose $w_{i}$ such that $\max\{w_{L},w_{H}\}<w_{i}$, the optimal choice for $w_{i}$ is $w_{H}+\kappa$, where $\kappa$ is the minimum weight increment as specified by the competition rules, given that $q(w_{i})$ is a decreasing function of $w_{i}$. Second, if lifter $i$ chooses $w_{i}$ such that $w_{i}<\min\{w_{L},w_{H}\}$, the optimal choice is the $w_{i}$ that maximizes $q(w_{i})P(w_{i}>y_{H})$, which leads to $w_{i}=y_{H}+\kappa$. Third, if lifter $i$ selects $w_{i}$ such that $w_{L}<w_{i}<w_{H}$, the optimal choice is $w_{i}$ that maximizes $q(w_{i})(1-q(w_{H}))P(w_{i}>y_{H})$, which also results in $w_{i}=y_{H}+\kappa$. Further analysis to precisely determine the analytical optimal attempt weight $w_{i}$ across all domains of attempt weights would require additional specifications for $q(\cdot),q(w_{H}),q(w_{L}),\lambda$, and other parameters, which increases the complexity of the modeling, identification, and estimation processes, and is beyond the scope of this study and our data.

### v4

@@ 8 Conclusion
Counterfactual simulations show that without competitive pressure, lifters tend to adopt more conservative attempts. However, the effects on performance are highly heterogeneous: while some lifters benefit from the absence of pressure, others experience lower success rates and reduced expected outcomes. These results highlight that external competition plays a dual role—motivating some lifters to perform better while destabilizing others—underscoring the importance of individual sensitivity in shaping performance under pressure.
@@
Appendix A Additional Results
A.1 Alternative definitions of competitive pressure
As a further robustness check, we consider alternative definitions of the pressure variables. In our baseline specification, pressure from a lower-ranked rival is measured as the difference between the lower rival’s declared attempt weight and the focal lifter’s current best outcome, while pressure from a higher-ranked rival uses the predicted attempt weight of the higher rival. Here, we replace these with two alternatives: (i) using the rival’s personal best record in place of the rival’s declared or predicted attempt weight, and (ii) using the rival’s best realized outcome in attempts preceding the focal attempt. Column (1) in Tables 14 and 15 re-estimates the baseline specification on the harmonized sample used in Columns (2) and (3), where rival personal-best and rival best-realized-outcome measures are available. Column (2) uses the rival’s personal best, and Column (3) uses the rival’s best realized outcome.
Table 14: Robustness: Alternative Pressure Definitions, Second Attempt
(1) (2) (3)
Dependent Variable $W_{it}^{2}$-(best) $W_{it}^{2}$-(best) $W_{it}^{2}$-(best)
Male -1.254*** 6.860*** 3.129***
(0.460) (1.222) (1.042)
Body weight 0.267*** 0.299*** 0.539***
(0.025) (0.060) (0.049)
Num experience 0.166 0.057 0.176
(0.115) (0.169) (0.144)
1(first participation) 31.479*** 36.194*** 33.560***
(1.372) (1.719) (1.705)
Pressure, lower rival (original) 0.101***
Pressure, higher rival (original) 0.425***
Pressure, lower rival (rival PB) 0.098***
Pressure, higher rival (rival PB) 0.127***
Pressure, lower rival (rival best outcome) 0.125***
Pressure, higher rival (rival best outcome) 0.323***
Control X X X
Num.Obs. 204215 204242 204242
R2 0.514 0.370 0.375
R2 Adj. 0.512 0.368 0.373
Notes: We control for gender, body weight, equipment category, age class, division, weight class, federation, and indicators for tie in declared weights and mismatch between lifting order and interim ranking. Column (1) uses the baseline pressure definition, where pressure from a lower-ranked rival is the difference between the lower rival’s declared attempt weight and the focal lifter’s current best outcome, and pressure from a higher-ranked rival is the difference between the higher rival’s predicted attempt weight and the focal lifter’s current best outcome. Column (2) replaces the rival’s declared or predicted attempt weight with the rival’s personal best. Column (3) replaces it with the rival’s best realized outcome in attempts preceding the focal attempt. Column (1) is estimated on the same restricted sample as Columns (2) and (3) for comparability, which yields fewer observations than in the full baseline table because rival personal-best information must be available. $^{*}p<0.1$; $^{**}p<0.05$; $^{***}p<0.01$.
Table 15: Robustness: Alternative Pressure Definitions, Third Attempt
(1) (2) (3)
Dependent Variable $W_{it}^{3}$-(best) $W_{it}^{3}$-(best) $W_{it}^{3}$-(best)
Male 3.439*** 8.984*** 5.175***
(0.633) (1.339) (1.023)
Body weight 0.347*** 0.316*** 0.500***
(0.049) (0.063) (0.064)
Num experience 0.178 0.069 0.171
(0.149) (0.192) (0.164)
1(first participation) 35.362*** 39.355*** 36.833***
(1.821) (2.046) (1.999)
Pressure, lower rival (original) 0.071***
Pressure, higher rival (original) 0.415***
Pressure, lower rival (rival PB) 0.056***
Pressure, higher rival (rival PB) 0.090***
Pressure, lower rival (rival best outcome) 0.033***
1 (0.012)
Pressure, higher rival (rival best outcome) 0.381***
Control X X X
Num.Obs. 181232 181232 181232
R2 0.369 0.273 0.307
R2 Adj. 0.366 0.270 0.304
Notes: We control for gender, body weight, equipment category, age class, division, weight class, federation, and indicators for tie in declared weights and mismatch between lifting order and interim ranking. Column (1) uses the baseline pressure definition, where pressure from a lower-ranked rival is the difference between the lower rival’s declared attempt weight and the focal lifter’s current best outcome, and pressure from a higher-ranked rival is the difference between the higher rival’s predicted attempt weight and the focal lifter’s current best outcome. Column (2) replaces the rival’s declared or predicted attempt weight with the rival’s personal best. Column (3) replaces it with the rival’s best realized outcome in attempts preceding the focal attempt. Column (1) is estimated on the same restricted sample as Columns (2) and (3) for comparability, which yields fewer observations than in the full baseline table because rival personal-best information must be available. $^{*}p<0.1$; $^{**}p<0.05$; $^{***}p<0.01$.
Across all specifications, the pressure coefficients remain positive and statistically significant at the 1% level for both the second and third attempts. The magnitudes vary somewhat across definitions, reflecting differences in the information content of each measure. The baseline specification, which uses declared or predicted attempt weights, yields the highest $R^{2}$, consistent with the fact that actual strategic choices incorporate more competition-relevant information than historical records alone. Nevertheless, the qualitative conclusions are unchanged: competitive pressure from both lower- and higher-ranked rivals significantly increases attempt weight selection.
A.2 First attempt weight and rival personal best
We examine whether first-attempt weight selection is associated with the personal bests of adjacent rivals. In our main specification, competitive pressure is defined with respect to potential changes in the interim ranking, which serves as the reference point in the sequential competition structure. Because no interim ranking exists prior to the first attempt, our baseline pressure variables cannot be defined for this stage. For this appendix exercise, we fix the order using declared first-attempt weights (with the official lifting order), define adjacent lower and higher rivals within that fixed order, and then construct pressure variables from personal-best gaps relative to those rivals. Unlike our main specification, rivals’ personal bests are determined prior to the competition and do not arise from the sequential realization of outcomes during the competition. As a result, they do not capture the exogenous variation generated by the sequential competition structure, and the results should therefore be interpreted as descriptive rather than causal.
Specifically, pressure from a lower rival is defined as (adjacent lower rival’s personal best) - (focal lifter’s personal best), and pressure from a higher rival is defined analogously. Table 16 reports the results. The dependent variable is $W_{it}^{1}-\text{(best)}$, the difference between the first-attempt weight and the lifter’s personal best. The coefficients on the rival personal best pressure variables are $0.111$ (lower rival) and $0.071$ (higher rival), both statistically significant at the 1% level. The positive signs are consistent with the pattern observed in our main second- and third-attempt results: lifters who face stronger rivals—as measured by the gap in personal bests—tend to select higher attempt weights. However, because personal bests are pre-determined and not generated by the within-competition information revelation mechanism that our identification strategy exploits, this finding should be interpreted as a descriptive association rather than a causal effect.
Table 16: First Attempt Weight and Rival Personal Best
Dependent Variable $W_{it}^{1}$-(best)
Male 22.713***
Body weight 0.746***
Num experience 1.090***
1(first participation) 128.328***
Pressure, lower rival (PB) 0.111***
Pressure, higher rival (PB) 0.071***
Control X
Num.Obs. 267342
R2 0.856
R2 Adj. 0.855
Notes: The dependent variable is the difference between the first-attempt weight and the lifter’s personal best ($W_{it}^{1}-\text{(best)}$). Rivals are defined as adjacent competitors in the fixed order based on declared first-attempt weights (using official lifting order when ties occur). Pressure from a lower-ranked rival is the difference between the lower rival’s personal best and the focal lifter’s personal best; pressure from a higher-ranked rival is defined analogously. We control for gender, body weight, equipment category, age class, division, weight class, federation, number of competition experiences, and an indicator for first-time participation. $^{*}p<0.1$; $^{**}p<0.05$; $^{***}p<0.01$.